OpenStudy (anonymous):

Determine whether the series is convergent or divergent. (1/3)+(1/6)+(1/9)+(1/12)+(1/15).... I figured out that a(subn) is (1/3)(1/n) but dont know where to go from there...

OpenStudy (anonymous):

\[\frac{1}{3}+\frac{1}{6}+\frac{1}{9}+\frac{1}{12}+\frac{1}{15}+\cdots=\sum_{n=1}^\infty \frac{1}{3n}\] so you're right about the series representation. To show convergence/divergence, there are a few tests you can use. Do you know the \(p\)-series test? Or the integral test?

OpenStudy (anonymous):

i only learned the test for divergence...if it doesnt = 0 then its def divergent but if it does - 0, then i have to take the limit of it

OpenStudy (anonymous):

does = 0*

OpenStudy (anonymous):

i meant if the limit doesnt = 0, then its def divergent, but if it does, then i have to take the limit and see what i get

OpenStudy (anonymous):

my book says that its actually divergent and am trying to figure out why

OpenStudy (anonymous):

I was thinking of sequence, sorry about that.

OpenStudy (anonymous):

its cool, no worries.

OpenStudy (anonymous):

You will use the p-series test on this one.

OpenStudy (anonymous):

http://www.youtube.com/watch?v=gB43EA9659c

OpenStudy (anonymous):

This guy explains it well

OpenStudy (anonymous):

im not familiar with that test. I just learned this stuff today and the only test i learned is the test for divergence where if the limit as n approaches infinity doesnt = 0 then its def divergent, but if it does, then i have to take the limit

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You have to find the limit of the sequence of partial sums: \[\left\{\frac{1}{3},\frac{1}{3}+\frac{1}{6},\frac{1}{3}+\frac{1}{6}+\frac{1}{9},\cdots\right\}_{n=1}^\infty\\ \left\{\frac{12}{36},\frac{18}{36},\frac{22}{36},\frac{25}{36},\cdots\right\}_{n=1}^\infty\] I can't remember the procedure for finding a formula for the \(n\)th partial sum, unfortunately... But it looks like each successive term increases without bound, and so the series diverges.

OpenStudy (anonymous):

ok thanks siths

OpenStudy (anonymous):